OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form of ax^2+bx+c=0 solution:2, only solution

OpenStudy (anonymous):

Single solution means in factored form you have\[(x-2)^2=0\]expand the binomial\[x^2-4x+4=0\]

OpenStudy (anonymous):

(x-2)^2=0 x^2-4x+4=0